OpenStudy (anonymous):

Can someone please check my coordinates and tell me If this counts as a positive slope A=5,2 B=5,6 C=2,18 D=8,12 E=9,7 F=8,1 G=12,12 H=14,14 I=18,7 J=20,6 I picked 8,12 and 12,12 slope=y2-y1/x2-x1 12-12/12-8 = 0/4 so would that be a positive slope... i'm not very good at geometry I just would like to check my work thank you very much for the help

OpenStudy (the_fizicx99):

What're you looking for exactly, pick 2 points from A - J and find the slope, put it into y = mx + b and see if it's a positive slope?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I'm not sure if im exactly doing it correctly because i'm not the best at geometry but i'm trying (:

OpenStudy (anonymous):

but I think instead of y=mx+b it's slope=y2-y1/x2=x1

OpenStudy (the_fizicx99):

K, um, true the slope formula is M = y2-y1/x2-x1 You picked D and G, correct? Umm, 0/4 is equal to 0. 0 is a horizontal line. Let's try to find another pair?

OpenStudy (anonymous):

Okay I will try another pair what about 9,7 and 8,1 I think the slope is 6

OpenStudy (anonymous):

Slope=y2-y1/x2-x1 so 1-7/8-9 = -6/1 = 6(: am I correct?

OpenStudy (the_fizicx99):

Let's use F and J, \(\ F = (8,1) - J = (20,6) \) \(\ M = \dfrac{y2 - y1}{x2-x1} \) \(\ M = \dfrac{6 - 1}{20-8} = \dfrac{5}{12} \) y-y1 = m(x-x1) y-1= 5/12(x-8) y - 1 = \(\ \dfrac{5}{12}x - \dfrac{10}{3} \) +1 +1 y = \(\ \dfrac{5}{12}x - \dfrac{7}{3} \) Now plot the y intercept at (0,-2.33) and apply rise / run. You should get something like this: http://prntscr.com/34q7m6 I did my calculations on wolfram alpha to avoid going to the fraction process >.> you can check if you'd like.

OpenStudy (the_fizicx99):

Remember that a positive slope is something like this, 

OpenStudy (anonymous):

Okay I think I understand what you did that there makes a lot more sense than what I was doing I just thought you had to come out with a positive number.. I always get confused when It goes to a decimal I overthink it and make it harder so that's what it would look like positive and those coordinates together make a positive slope? awesome(: thank you. I also need to find Negative slope do you think I could try that out and you tell me If i'm correct as well? I would appreciate it so much

OpenStudy (the_fizicx99):

Sure, yes a positive slope is kinda like; / -- lol think of it as climbing up a mountain xD a negative slope is like going down a mountain :p

OpenStudy (anonymous):

Awesome haha thank you so just give me a few minutes to work it out please(: thanks!

OpenStudy (the_fizicx99):

The y intercept can be positive or negative. What's important is the slope, if it's positive it'll always be a positive slope because it's a constant rise/run

OpenStudy (the_fizicx99):

Sure

OpenStudy (anonymous):

Okay so I'm a bit confused but I think i'm doing it right I did 18,7 and 20,6 the slope is -0.5 and in the graph it looks like the negative mountain thing you were talking about did I do that correctly (: @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Which letters did you chose :o

OpenStudy (anonymous):

I and J

OpenStudy (the_fizicx99):

I wouldn't use J again, your teacher might want you to use another, but it's ok I suppose if she's isn't annoying cx (18,7) & (20,6) m = 6 - 7 / 20 - 18 = -1 / 2 ... keep -1/2 so that you can apply rise/run Yeah, the slope is correct, use y-y1=m(x-1)

jigglypuff314 (jigglypuff314):

I suggest reordering them from lowest x to highest x then taking the slope of the coordinate that has the lowest x and the coordinate that has the highest x... http://www.wolframalpha.com/input/?i=%285%2C2%29+%285%2C6%29+%282%2C18%29+%288%2C12%29+%289%2C7%29+%288%2C1%29+%2812%2C12%29+%2814%2C14%29+%2818%2C7%29+%2820%2C6%29

OpenStudy (the_fizicx99):

Jiggy that's a bit too much if she doesn't understand it >.>

OpenStudy (anonymous):

This helped me so much I appreciate it so much(: I'm surprised I actually did that right lol this will help me with my next question it's asking me which coordinates are parallel and that has it graphed and everything you guys are great!

OpenStudy (the_fizicx99):

Yw :>

OpenStudy (anonymous):

@jigglypuff314 @tHe_FiZiCx99 I'm not sure if you guys are still on but can you please check my work on one more thing please(: I will post the work below

OpenStudy (anonymous):

For this one it's using the same coordinates I just need to find which points create a Perpendicular line I chose H and I so y-7=(-1/14) times (x-18) m=-1/14 times x1=18 and y=7 y-7 (-1/14 times x+ (1/14)(18) distribute -1/14 y-7=(-1/14 times x+18/14) y=(-1/14) times x+18/4+7 add 7 to isolate y y=-1/14x+58/7 reduce the equation of the line that is perpendicular to y=14 times x+14 and goes through (18,7) is y=(-1/14 x+58/7) is that correct? i'm sorry I hope this isn't inconvience

OpenStudy (anonymous):

I thought about what you said about not using the same coordinate sover again so I feel like maybe I should try to find a different pair that makes a perpendicular line just to be safe

OpenStudy (the_fizicx99):

I believe you need to have 2 equations to have a perp line lol, I think you misunderstood the y-y1=m(x+x1) >.< hm xD, basically just find 2 pairs, find the slope, put it in y-y1=(x-x1) and plot it :> Do the same thing over, should look something like:  it's a 90 degree >.> looks horrible ;-;

OpenStudy (anonymous):

I thought that was what I did though :o ugh i'm sorry I am a bit confused. Okay i'm going to choose different points this time

OpenStudy (the_fizicx99):

kk, I have to get off D; doing my own project atm :(

OpenStudy (anonymous):

Okay I understand thank you so much for the help (: I will try to do my best on the rest. Good luck on your project!

OpenStudy (the_fizicx99):

Thanks xD I'm sure @jigglypuff314 can help you Cx

jigglypuff314 (jigglypuff314):

psh wut? xD I can try, but idk whut going on :P

OpenStudy (anonymous):

Haha I have to find out of my coordinates I posted above what would make a perpendicular line.. I'm just a bit confused but i'm trying to figure it out

OpenStudy (anonymous):

so I choose coordinates and then using their perpendicular formula separately I work them together to see if it's perpendicular? @jigglypuff314

jigglypuff314 (jigglypuff314):

try making equation out of (12, 12) and (14, 14) then perpendicular might be (12, 12) to (20, 6)

OpenStudy (anonymous):

Okay so first do the coordinates (12,12) (14,14) and using that perpendicular line formula I then use it with the other coordineates formula to figure out if they intersect

jigglypuff314 (jigglypuff314):

yeah sure >.> I'm not quite sure what your question was, but okay xD

OpenStudy (anonymous):

lol it's okay I am still really confused and don't really get what I'm doing I think I will contact my teacher about it because I don't want to be a inconvenience to you thank you though

jigglypuff314 (jigglypuff314):

it's fine ^_^ perpendicular just means that slopes are flipped opposites to each other

OpenStudy (anonymous):

so I plug the slopes in to see if it graphs as perpendicular... I sound so stupid I really just dont understand it but seriously thanks for trying to explain